# Enhancing LLM Reasoning via Non-Human-Like Reasoning Path Preference Optimization

### v1

@@ 5 Conclusion
In this paper, we propose CGPO, a DPO-based step-wise reasoning path preference optimization method. Our method emphasizes non-human-like reasoning paths in two key ways: 1) the training data does not rely on strong models; 2) the split for reasoning paths does not depend on predefined anchors. We show the effectiveness of CGPOin both mathematics and code domains, achieving significant improvements and generalization ability over approaches that use data generated by strong models. Furthermore, our experiments reveal that, within a certain range, intermediate reasoning path-focused optimization leads to greater performance gains than whole reasoning path optimization.
Limitations
Due to computational constraints, we are unable to conduct experiments on larger models (e.g.,70B) to explore the scalability with respect to model size. In addition, our experiments were limited to the code and math domains, primarily because fine-grained evaluation in the commonsense domain is challenging. However, we believe our approach remains applicable as long as a general fine-grained scoring model is available, since the proposed methodology is domain-agnostic.
@@ A.2 Statistics of the Positional Relationship
Table 6: Statistics on the relative positions of first error step and lowest confidence point, where "Before" and "After" denote the first error before or after the model confusion point
@@
"Analyze the following text and identify the first location where an error occurs, reporting the exact substring/token, its position (line and character index), why it is an error, and why this is the first error (i.e., why all preceding text is error-free); if no error exists, reply ‘No error found’; Text:" [Q-A pair]

### v2

@@ 5 Conclusion
In this paper, we propose CGPO, a DPO-based step-wise reasoning path preference optimization method. Our method emphasizes non-human-like reasoning paths in two key ways: 1) the training data does not rely on strong models; 2) the split for reasoning paths does not depend on predefined anchors. We show the effectiveness of CGPO in both mathematics and code domains, achieving significant improvements and generalization ability over approaches that use data generated by strong models. Our analyses show that exploring non-human-like reasoning paths enhances model capabilities and is more effective than enforcing human-like reasoning paths. Separately, our experiments reveal that, within a certain range, intermediate reasoning path-focused optimization leads to greater performance gains than whole reasoning path optimization.
Impact Statements
CGPO is an automated, scalable, and general method designed to optimize reasoning trajectories in LLMs through confidence-guided preference learning. By shifting away from labor-intensive human annotations, CGPO enables the autonomous generation and selection of high-quality reasoning paths, making it readily adaptable to a vast array of complex tasks. Our investigation into non-human-like, model-native reasoning paths offers a new perspective on optimization algorithms.
@@ A.2 Statistics of the Positional Relationship
Table 5: Statistics on the relative positions of first error step and lowest confidence point, where ”Before” and ”After” denote the first error before or after the model confusion point
@@
”Analyze the following text and identify the first location where an error occurs, reporting the exact substring/token, its position (line and character index), why it is an error, and why this is the first error (i.e., why all preceding text is error-free); if no error exists, reply ‘No error found’; Text:” [Q-A pair]
@@
A.3 Training Details
The key hyperparameters we used in training the model are presented in Table 6, including the learning rate, beta, batch size, and number of epochs. For all models with fewer than 32B parameters, we set $\beta$ to 0.1, while for 32B model, we set $\beta$ to 0.3.
Table 6: Training hyperparameters for all models.
Model $\beta$ LR Batch Size Epochs
MetaMath-Mistral-7B 0.1 $5\times 10^{-7}$ 128 4
MetaMath-Llama-8B 0.1 $5\times 10^{-7}$ 128 4
Qwen2-7B-SFT 0.1 $5\times 10^{-7}$ 128 4
DeepSeekMath-Base-7B-SFT 0.1 $5\times 10^{-7}$ 128 4
Qwen1.5-32B-SFT 0.3 $5\times 10^{-7}$ 128 4
DeepSeek-R1-Distill-Qwen-1.5B 0.1 $5\times 10^{-7}$ 128 4
DeepSeek-R1-Distill-Qwen-7B 0.1 $5\times 10^{-7}$ 128 4
Deepseek-Coder-7B-Instruct-v1.5 0.1 $5\times 10^{-7}$ 128 4
A.4 CGPO performance on different sampling temperatures
We experimented with using different temperatures in CGPO to sample non-human-like data. The figure 7 show that, across all evaluated temperatures, CGPO consistently outperforms the base model, demonstrating its robustness to the sampling temperatures.
Figure 7: The relationship between the sampling temperatures and the trained model performance.
A.5 CGPO Training Data Analysis
Table 7: Samples and Proportion of decision tokens for CGPO training dataset.
Category Lowest-conf. (%) All Tokens (%) Sample Snippet
Math Reasoning
    Number 17.99 17.45 sum is an integer, the answer is 2000.
    Proper Noun 8.30 4.44 7000 miles: Bangkok-Honolulu
    Symbol 9.45 8.58 90 + 90 + … + 71 = 420 degrees.
Logical Relation
    Conjunction 9.15 3.97 We then add and subtract \$25\$
    Verb 9.10 7.74 we calculate the sum of 46 and 37
Supportive Elements
    Particle 4.29 1.51 positive integers that do not have
    Auxiliary Verb 0.50 5.97 we have $y^{2}+12y=45$
Others 41.22 50.34
We categorize 21,374 lowest-confidence tokens using a rule-based lexical taxonomy into three categories: Math Reasoning, Logical Relation, and Supportive Elements. Table 7 presents the distribution of these decision-related token categories in the CGPO training dataset compared to their overall frequency in all tokens. We observe that tokens associated with reasoning and computation are significantly enriched at the model’s lowest-confidence points. Specifically, numeric tokens, computation symbols, and computation-related verbs and prepositions occur more frequently among decision tokens than in the overall token distribution. This enrichment indicates that the lowest-confidence points selected by CGPO are strongly correlated with reasoning-related operations.